OpenStudy (anonymous):

4(-3m^3n^4p)^2/(9mn^2p^9)

OpenStudy (anonymous):

I am getting very confused.

hero (hero):

Easy! lol

OpenStudy (anonymous):

lol then can you help?

OpenStudy (anonymous):

hello? anyone there..

OpenStudy (anonymous):

take your time...

OpenStudy (anonymous):

what do you need?

OpenStudy (anonymous):

\[\frac{4(-3m^3n^4p)^2}{9mn^2p^9}\]is the problem?

OpenStudy (anonymous):

\[\frac{4(-3m^3n^4p)^2}{9mn^2p^9}=\frac{4\times (-3)^2m^6n^4p^2}{9mn^2p^9}\]i

OpenStudy (anonymous):

now cancel to get \[\frac{4m^5n^2}{p^7}\]

OpenStudy (anonymous):

Thanks so much I get it now!

OpenStudy (anonymous):

I was confused and thought I had to multiply the four first but I now know I needed to use the orders of operation